the measure of the two consecutive angles of a parallelogram are 3x + 45 and 6x respectively. find the measures of the four angles.
Answer by htmentor(1343)

Consecutive angles are supplementary, so we can write
3x+45 + 6x = 180
Solve for x:
9x = 135
x = 15
So the two angles are 6*15 = 90, 3*15+45 = 90
The opposite angles are congruent, so all 4 angles are 90 deg, i.e. a square.